Subject: APO 3.5m users committee, minutes for 1/23/12 meeting

From: Suzanne Hawley

Submitted: Wed, 1 Feb 2012 21:50:06 -0800 (PST)

Message number: 1260 (previous: 1259, next: 1261 up: Index)

 		APO 3.5-m Users Committee Phonecon, 1/23/12

Attending:  Suzanne Hawley, Mark Klaene, Bill Ketzeback, Al Harper, Don 
York, Remy Indebetouw, Jon Holtzman, John Bally, Michael Strauss, Scott 
Anderson, and Bruce Gillespie

Absent:  Sean Moran


User feedback and comments from institutional representatives:

o Chicago (Don York) - Don said that everything is fine.

o Washington (Scott Anderson) - Scott had nothing to report.

o New Mexico State (Jon Holtzman) - Jon had nothing to report.

o Princeton (Michael Strauss) - Michael mentioned two items: the 
TripleSpec pattern noise (discussed below), and a PU user had difficulty 
working with the site staff during a spate of NIC-FPS image readout 

o Colorado (John Bally) - John mentioned two items: the future of 
narrow-band imaging at APO is a concern (see writeup in last months 
minutes), and whether NIC-FPS will be sent to UVa for repairs.

o Virginia (Remy Indebetouw) - Remy had no user issues.  He mentioned 
financial concerns that the UVa faculty have regarding the impending loss 
of ARC partners.  Suzanne said that ARC is actively pursuing new partners, 
and suggestions are welcome.

o Johns Hopkins (Sean Moran) - No report.


Discussion of telescope/instruments report:

Mark said that DIS, Agile, SPIcam, and the telescope worked reliably 
during the period.  There is noticeable pattern noise on TripleSpec frames 
that we are working with UVa to diagnose.  The noise appears to be a 19 
kHz signal, and may be related to a power supply.  The noise is reduced 
with FS8, and has been seen at a much lower level since April 2010.

The new echelle grating is en route to APO.  JHU engineers visited APO in 
December to evaluate the mechanics of mounting the grating, and everything 
looks OK.  The grating swap is planned for May/June.  The new echelle 
slitviewer camera system has arrived, and is undergoing tests.

The NIC-FPS image-corruption problem still occasionally happens, but is 
not seriously affecting science.  The GFP IFU commissioning continued, and 
it may be integrated with TUI in February.

Suzanne added that in the realm of new instrumentation for the telescope, 
there are many trade-space options to consider between GRIZLI, GFP, 
SPIcam, and NIC-FPS that bear on the use of floor space, wide-field 
imaging, and narrow-band imaging.  These trades will continue to be a 
discussion topic at these meetings.

Mark Klaene's detailed report for the reporting period follows:


 	      3.5-m Telescope and Instruments Highlights
 	                 12/7/11 through 01/16/12

0) Overview

We have moved more into the forecasted la Nina weather pattern with less 
frequent and dryer storms.  Systems have continued to move through the 
area about 1/week but with little precipitation.  Clouds and humidity have 
hampered observing at times.

Telescope continued to run well and with few instrument problems.  Most of 
the staff were able to avoid having to come in over the holiday break and 
still meet the observing schedule.

1) Telescope

The telescope has been operational with only minor problems reported. 
With lighter snowfall, clearing of the enclosure roof has not been a major 
headache.  Icing on the lightning arrestors has caused a couple of nights 
with stricter wind-closure criteria.

2) Instruments

DIS replacement Blue 400 grating is still at JHU awaiting testing. 
Expect it on the mountain by early February if it is okay.

Agile and SPIcam operated without issues.

Some TSpec images have showed a pattern noise at a low level.  Joe 
Huehnerhoff has information regarding these images and their frequency of 
occurrence at  UVa and 
APO staff are working on identifying the source.  At this point the likely 
suspect is a power supply.

Echelle is operational.  JHU personnel have reviewed the grating 
installation and a new grating has been purchased and received.  Plans for 
installation are being developed.  We expect this will occur in May or 
June of this year.  The new guide camera has arrived on site.  Bench 
testing and software development will begin over the next few weeks.

NIC-FPS has been operating in shared-risk observing status.  The image 
corruption issues are still present on occasion but have not been reported 
to be grossly affecting science observations.  Recently the problem did 
cause a fair amount of lost time and observing inefficiency.  A reported 
vacuum problem over the holiday break appears to have worked itself out.

GFP software development is continuing well, a light-leak fix is in work. 
Commissioning has slowed due to funding issues at GSFC.


ToO feedback:

Suzanne reported that we have recently handled numerous ToO events at the 
telescope, and asked if there was any feedback to discuss.  Michael 
Strauss thanked everyone for cooperating with the Princeton PI who is 
invoking many of the ToOs.  The issue of a ToO program being organized, 
and then the PI deciding not to observe the ToO at the last minute was 
brought up.  Suzanne said that this is awkward, because it can be 
difficult to contact all the affected parties, who may have made other 
plans. Thus ToO programs should come to the telescope with a backup 
program so that they can use the time profitably. Don York mentioned that 
a recent APO ToO bulletin was released without it being shown to all the 
authors.  Suzanne reminded everyone that the policy on notices and 
bulletins is that they be shown to all authors before publication, 
including the PI of the program that was pre-empted, who should be invited 
to be an author (but who is free to decline if they wish). Further 
feedback about ToO programs is sought, especially as their frequency 
increases.  In particular, Suzanne wants to know if the payback scheme is 
proving useful to the PIs of the pre-empted programs.


NIIS in March:

Mark announced the upcoming arrival of the UWyoming NIIS visiting 
instrument at the end of February.  It is a 12 arcminute FOV NIR camera 
that will be on the telescope during the first week of March.  ARC users 
are invited to submit observing targets for this trial run; Bill Ketzeback 
is the contact person (  We would especially 
appreciate users suggesting targets that will help characterize the 
instrument, e.g. using the full FOV.  Users interested in learning more 
about NIIS can go to the UWyoming website for information:

and can also look at the lab performance white paper:


GFP in Q2:

Suzanne said that if there are potential users of GFP in the second 
quarter, they should contact Bruce Woodgate ( 
to discuss its status.


Plans for coming year:

Suzanne will be on sabbatical during July 2012-June 2013.  One of the jobs 
she does as Director is to manage the quarterly schedule, with Russet's 
help.  This includes ToO events, open time, unexpected 
engineering/maintenance, time trades, etc..  She would like to engage an 
ARC astronomer to volunteer to help with the telescope scheduling, 
particularly during periods when she is unavailable during her sabbatical. 
Interested parties should contact her directly.

Suzanne announced that Michael Blanton (NYU) is the Director-designate for 
AS3 (the after SDSS-III project currently being planned for the 2.5m 
telescope at APO).  She encourages people to get involved with AS3 as it 
starts ramping up this year.



[open from previous meeting]:


[new actions from this meeting]:



New business:

Suzanne said that we are planning to put a new page on the 3.5-m 
instrument web site that list current instrument status and news. The 
location will be sent to the users when it is up and running, and users 
will be encouraged to check this site regularly when planning for 
observations and prior to observing.

Note added after meeting:

The instrument status website is now available, see:


Next meeting:  The next Users Committee phonecon will be on Monday 27 
February at 8:30 AM Pacific Time.  The agenda and other materials will be 
sent to the committee members during the preceding week.

APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 1260 in the apo35-general archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to